# Webs Generated by Products of convex and homogeneous Foliations on $\mathbb{P}^{2}$

### v1

@@ 3. Homogeneous Foliations
Take $s\in\Sigma_{\mathcal{H}_{1}}^{rad}\setminus(\Sigma_{\mathcal{H}_{2}}^{rad}\cup\Sigma_{\mathcal{H}_{3}}^{rad})$, we have that $s$ is a radial singularity of order $\nu_{1}$ of $\mathcal{H}_{1}$ and given a generic line $l$ (particularly, not invariant) on $\check{s}$, we can write $\operatorname{Leg}\mathcal{H}_{1}=\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}$ with $\mathcal{W}_{\nu_{1}}$ leaving $\check{s}$ invariant and with $\Delta(\mathcal{W}_{\nu_{1}})$ having minimal multiplicity along $\check{s}$ and the $(d_{1}-\nu_{1})$-web $\mathcal{W}_{d_{1}-\nu_{1}}$ is transverse to $\check{s}$ As $s\not\in\Sigma_{\mathcal{H}_{2}}^{rad}\cup\Sigma_{\mathcal{H}_{3}}^{rad}$ we have that $\operatorname{Tang}(\mathcal{H}_{2},l,s)$ and $\operatorname{Tang}(\mathcal{H}_{3},l,s)$ are less or equal than $1$.
  • If $s\not\in(\operatorname{Sing}\mathcal{H}_{2}\cup\operatorname{Sing}\mathcal{H}_{3})$, we can assume $l\cap(\operatorname{Sing}\mathcal{H}_{2}\cup\operatorname{Sing}\mathcal{H}_{3})=\emptyset$, then $\operatorname{Leg}\mathcal{H}_{2}$ and $\operatorname{Leg}\mathcal{H}_{3}$ are completely transversal to $\check{s}$. On the other hand, for example, if one direction of $\mathcal{W}_{d_{1}-\nu_{1}}$ and $\operatorname{Leg}\mathcal{H}_{2}$ coincide, this direction would correspond to a point of $\operatorname{Tang}(\mathcal{H}_{1},\mathcal{H}_{2})=L_{\infty}\cup l_{1}\cup\cdots\cup l_{k}$ therefore, we would have $l=l_{j}$ for some $j$, which does not occur.Thus, $\operatorname{Leg}\mathcal{H}_{2}$ and $\operatorname{Leg}\mathcal{H}_{3}$ are transverse to $\mathcal{W}_{d-\nu_{1}}$ and in the same way they are mutually transversal, thus in the neighborhood of $l$ we can write $\operatorname{Leg}\mathcal{W}=\mathcal{W}_{\nu_{1}}\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{H}_{2}\boxtimes\operatorname{Leg}\mathcal{H}_{3})$ with $\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\operatorname{Leg}\mathcal{H}_{2}\boxtimes\operatorname{Leg}\mathcal{H}_{3}$ smooth and transverse to $\check{s}$. Therefore, by [6, Proposition 2.6], the curvature of $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$.
  • If $s\in(\operatorname{Sing}\mathcal{H}_{2}\setminus\operatorname{Sing}\mathcal{H}_{3})$ we have $\operatorname{Tang}(\mathcal{H}_{2},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{H}_{2}=\mathcal{F}_{2}\boxtimes\mathcal{W}_{d_{2}-1}$ where $\mathcal{F}_{2}$ is a foliation tangent to $\check{s}$ and $\mathcal{W}_{d_{2}-1}$ is a web transverse to $\check{s}$. By the previous argument, $\mathcal{W}_{d_{2}-1}$ and $\operatorname{Leg}\mathcal{H}_{3}$ are transverse to $\mathcal{W}_{d_{1}-\nu_{1}}$ and to each other as well. Thus, by [2, Proposition 3.9], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{F}_{2}\boxtimes\mathcal{W}_{\nu_{1}})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}\boxtimes\operatorname{Leg}\mathcal{H}_{3})$ is holomorphic along $\check{s}$.
  • If $s\in(\operatorname{Sing}\mathcal{H}_{2}\cap\operatorname{Sing}\mathcal{H}_{3})$ we have $\operatorname{Tang}(\mathcal{H}_{2},l,s)=\operatorname{Tang}(\mathcal{H}_{3},l,s)=1$ and we decompose $\operatorname{Leg}\mathcal{H}_{2}=\mathcal{F}_{2}\boxtimes\mathcal{W}_{d_{2}-1}$ and $\operatorname{Leg}\mathcal{H}_{3}=\mathcal{F}_{3}\boxtimes\mathcal{W}_{d_{3}-1}$ as before. By the previous argument, $\mathcal{W}_{d_{2}-1}$ and $\mathcal{W}_{d_{3}-1}$ are transverse to $\mathcal{W}_{d_{1}-\nu_{1}}$. Thus, by [2, Remark 3.10], the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{F}_{2}\boxtimes\mathcal{F}_{3}\boxtimes\mathcal{W}_{\nu_{1}})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-1}\boxtimes\mathcal{W}_{d_{3}-1})$ is holomorphic along $\check{s}$.
Now, take $s\in(\Sigma_{\mathcal{H}_{1}}^{rad}\cap\Sigma_{\mathcal{H}_{2}}^{rad})\setminus\Sigma_{\mathcal{H}_{3}}^{rad}$, we have that $s$ is a radial singularity of orders $\nu_{1}$ and $\nu_{2}$ of $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ respectively and we have local decompositions as before: $\operatorname{Leg}\mathcal{H}_{i}=\mathcal{W}_{\nu_{i}}\boxtimes\mathcal{W}_{d_{i}-\nu_{i}}$. We need to analyze two situations:
  • If $s\not\in\operatorname{Sing}\mathcal{H}_{3}$ by Lemma 1.3 and [6, Proposition 2.6], we obtain that the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}})\boxtimes(\operatorname{Leg}\mathcal{H}_{3}\boxtimes\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-\nu_{2}})$ is holomorphic along $\check{s}$
  • if $s\in\operatorname{Sing}\mathcal{H}_{3}$, we can decompose $\operatorname{Leg}\mathcal{H}_{3}=\mathcal{F}_{3}\boxtimes\mathcal{W}_{d_{3}-1}$ with $\mathcal{F}_{3}$ a foliation tangent to $\check{s}$. Again by Lemma 1.3 and [2, Remark 3.10] the curvature of $\operatorname{Leg}\mathcal{W}=(\mathcal{W}_{\nu_{1}}\boxtimes\mathcal{W}_{\nu_{2}}\boxtimes\mathcal{F}_{3})\boxtimes(\mathcal{W}_{d_{1}-\nu_{1}}\boxtimes\mathcal{W}_{d_{2}-\nu_{2}}\boxtimes\mathcal{W}_{d_{3}-1}$ is holomorphic along $\check{s}$.
In the case where $\check{s}$ is radial singularity of $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ of respective orders $\nu_{1}$, $\nu_{2}$, $\nu_{3}$ we obtain that $\operatorname{Leg}\mathcal{W}$ is holomorphic along $\check{s}$ by lemma 1.3 and [6, Proposition 2.6].
Then, $K(\operatorname{Leg}\mathcal{W})$ is holomorphic on ˇ$\check{\mathbb{P}}^{2}\setminus\check{O}$ and by Lemma 3.1 we obtain the flatness of $\operatorname{Leg}\mathcal{W}$.
@@
Let $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ be convex homogeneous foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of respective degrees $d_{1}$ and $d_{2}$, such that, $\operatorname{Tang}(\mathcal{H}_{1},\mathcal{H}_{2})=L_{\infty}\cup l_{1}\cup\cdots\cup l_{k}$ where $l_{j}$ is a line invariant $\forall j=1,\cdots,k$, and let $l$ be an invariant line. Consider the web $\mathcal{W}=l\boxtimes\mathcal{H}_{1}\boxtimes\mathcal{H}_{2}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}l\boxtimes\operatorname{Leg}\mathcal{H}_{1}\boxtimes\operatorname{Leg}\mathcal{H}_{2}$ is flat.
@@
    By Lemma 1.3 and [2, Proposition 3.9], we conclude that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic along $\check{a}$.
@@
To finish the proof of Theorem B we need some general results about curvature.
Proposition 3.5.
Let $\mathcal{W}=\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of webs. If for each $i,j,l=1,\cdots,n$ we have:
  1. i.
    $K(\mathcal{W}_{i})=0$;
  2. ii.
    $K(\mathcal{W}_{i}\boxtimes\mathcal{W}_{j})=0$, $i\neq j$;
  3. iii.
Then $K(\mathcal{W})=0$.
and
Corollary 3.6.
Let $\mathcal{W}=\mathcal{F}\boxtimes\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of a foliation and $n$ webs. If for each $i,j,l=1,\cdots,n$ the Proposition 3.5 hypotheses are satisfies and we have:
  1. i.
    $K(\mathcal{F}\boxtimes\mathcal{W}_{i})=0$;
  2. ii.
Then $K(\mathcal{W})=0$.
and
Now a simple computation shows that $K(\mathcal{W})=0$.
Corollary 3.7.
Let $\mathcal{W}=\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}\boxtimes\mathcal{W}_{1}\boxtimes\cdots\boxtimes\mathcal{W}_{n}$ be a product of two foliations and $n$ webs. With the same hypotheses as the previous lemma for $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, and additionally: if $K(\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}\boxtimes\mathcal{W}_{i})=0$ for each $i=1,\cdots,n$. Then $K(\mathcal{W})=0$.
Proof.:  With the same notation as before we have
Combining this with the information obtained in the previous lemmas, we conclude that $K(\mathcal{W})=0$.
Proof. of Theorem B: Set $\mathcal{W}^{\prime}=\operatorname{Leg}l_{i}\boxtimes\cdots\boxtimes\operatorname{Leg}l_{k}$ and $\mathcal{W}^{\prime\prime}=\operatorname{Leg}\mathcal{H}_{1}\boxtimes\cdots\boxtimes\operatorname{Leg}\mathcal{H}_{n}$. By [1, lemma 2.1], we have that
By [6, Theorem 4.2], Propositions 3.2, 3.3, and 3.4 we are in condition to apply Corollary 3.6 and conclude that $K(\operatorname{Leg}l_{i}\boxtimes\mathcal{W}^{\prime\prime})=0$. Similarly, we use in addition [1, Theorem 2] to be able to use Corollary 3.7 and obtain $K(\operatorname{Leg}l_{i}\boxtimes\operatorname{Leg}l_{j}\boxtimes\mathcal{W}^{\prime\prime})=0$. Finally, by Proposition 3.5 we have $K(\mathcal{W}^{\prime\prime})=0$ and we finish our proof.
@@
Let $\omega_{d}$ be the form $\omega_{d}=y^{d}dx-x^{d}dx$. It is shown in [3, Proposition 4.1] that this form defines a convex homogeneous foliation of degree $d$ with invariant lines: $L_{\infty}$, $x=0$, $y=0$, $y=\xi x$, with $\xi^{d-1}=1$.

### v2

@@ 3. Homogeneous Foliations
If $s\in(\Sigma_{\mathcal{H}_{1}}^{rad}\cup\Sigma_{\mathcal{H}_{2}}^{rad}\cup\Sigma_{\mathcal{H}_{3}}^{rad})$, we can proceed as in the Proposition 2.5 and obtain that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic over $\check{s}$. So since $K(\operatorname{Leg}\mathcal{W})$ is holomorphic on $\mathbb{P}^{2}_{\mathbb{C}}\setminus\check{O}$ by Lemma 3.1 we conclude that $\operatorname{Leg}\mathcal{W}$ is flat.
@@
Let $\mathcal{H}_{1}$, $\mathcal{H}_{2}$ be convex homogeneous foliations on $\mathbb{P}_{\mathbb{C}}^{2}$ of respective degrees $d_{1}$ and $d_{2}$, such that, $\operatorname{Tang}(\mathcal{H}_{1},\mathcal{H}_{2})=L_{\infty}\cup l_{1}\cup\ldots\cup l_{k}$ where $l_{j}$ is a line invariant $\forall j=1,\ldots,k$, and let $l$ be an invariant line. Consider the web $\mathcal{W}=l\boxtimes\mathcal{H}_{1}\boxtimes\mathcal{H}_{2}$. Then $\operatorname{Leg}\mathcal{W}=\operatorname{Leg}l\boxtimes\operatorname{Leg}\mathcal{H}_{1}\boxtimes\operatorname{Leg}\mathcal{H}_{2}$ is flat.
@@
    By [4, Theorem 1.1], we conclude that $K(\operatorname{Leg}\mathcal{W})$ is holomorphic along $\check{a}$.
@@
Proof. of Theorem B: With the results of this section added to those presented in the previous section, similarly to what was done in the demonstration of Theorem A, Theorem B is established.
@@
Let $\omega_{d}$ be the form $\omega_{d}=y^{d}dx-x^{d}dy$. It is shown in [3, Proposition 4.1] that this form defines a convex homogeneous foliation of degree $d$ with invariant lines: $L_{\infty}$, $x=0$, $y=0$, $y=\xi x$, with $\xi^{d-1}=1$.
@@
Example 3.6.
Let $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be foliations given by the respective $1$-forms $\omega_{1}=\lambda ydx+xdy$ and $\omega_{2}=y^{2}dx+x^{2}dy$. Consider the $3$-web $\mathcal{W}=\mathcal{F}_{1}\boxtimes\mathcal{F}_{2}$ and note that $\operatorname{Tang}(\mathcal{F}_{1},\mathcal{F}_{2})=L_{\infty}\cup l_{1}\cup l_{2}\cup T$ where the lines $l_{1}:\{x=0\}$ and $l_{2}:\{y=0\}$ are invariant by $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, and $T:\{y+\lambda x\}$ is transversal. Take $p=\mathcal{G}_{F_{1}}(T)$ then clearly $\check{p}\subset\Delta(\operatorname{Leg}\mathcal{W})$ and it is not difficult to verify, by barycenter criterion (see [4, Theorem 1.1]), that the curvature of $\operatorname{Leg}\mathcal{W}$ is not holomorphic in $\check{p}$.